two cars leave the same place at the same time and going to the same place
one is traveling at 43 km per hour
the other at 52 km per hour
how far apart will they be after 40 minutes

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

two cars leave the same place at the same time and going to the same place
one is traveling at 43 km per hour
the other at 52 km per hour
how far apart will they be after 40 minutes
-----
40 min = 2/3 hour
-------------------------
distance = rate*time
slower car distance:: 43(2/3) km
faster car distance:: 52(2/3)km
----
Distance apart = (52-43)(2/3) = 9(2/3) = 6 km
-------------
Cheers,
Stan H.
------------

Answer by ikleyn(52915)

.
two cars leave the same place at the same time and going "to the same place".
                                                          -----------------

VERY STRANGE CONDITION.

The desire to solve disappears.